OpenStudy (anonymous):

Solve by factoring: 3x^2 + 2x - 1 = 0 {-1, 2/3} {1, -2/3} {-1, 1/3} {1, -1/3}

OpenStudy (anonymous):

please explain how i find this

OpenStudy (jamesj):

well, dividing first by 3, we have \[ x^2 + \frac{2}{3}x - \frac{1}{3} = 0 \] Now, you want factors of -1/3 that add up to 2/3

hero (hero):

@EmmaH, the steps for solving this are the same as what I showed you in the previous problem.

OpenStudy (anonymous):

divide what by three?

hero (hero):

I'll even give you a hint, but you have to solve the rest yourself: Given: \(3x^2 + 2x - 1\) Split the middle term: \(3x^2 + 3x - 1x - 1\) Now factor by grouping

OpenStudy (anonymous):

(3x^2+3x) + (-1x-1)?

hero (hero):

Okay, now what's the next step?

hero (hero):

Do you know how to factor \(3x^2 + 3x\) ?

OpenStudy (anonymous):

3x(1x+1)

hero (hero):

Yes, or simply 3x(x + 1) Writing 1 before the x is not necessary. Now factor (-1x - 1). Let me know what you get.

OpenStudy (anonymous):

?

OpenStudy (jamesj):

Well, you have a factor of (x+1) above where you wrote 3x(x+1) Can you get a factor of (x+1) out of (-1x - 1) ?

OpenStudy (anonymous):

what do you mean?